You are given a 0.25 M SO32- solution some of which you place in a flask and a 0.25 M solution of the strong acid hydrochloric acid (HCl) which you place in a burette. You also place a pH probe in the flask to record the pH values. You set about titrating the SO32- with HCl.
If at the middle of a buffer region the pH is 7.25, what is a Kb value for SO32-?
Answer #1

In the middle of the buffer region the concentration of the weak acid ann its conjugate base is the same so:

pH=pKa+log\frac{[A^-]}{[HA]}

The second part of the equation is 0, and pKa=pH

Now, you want to calculate the Kb, and you now that pKa+pKb=14

So: pKb=14-pKa=14-7.25=6.75

and K_b=10^{-6.75}=1.78\times 10^{-7}